Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Manhattan Math: Probability

Expert replies
by Ashishkapoor7 » Thu Jul 19, 2012 5:16 am
Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

8/33
62/165
17/33
103/165
25/33
Join the discussion
Source: — Problem Solving |

by brainturner » Thu Jul 19, 2012 5:51 am
Bill has 12 cards in his deck. After shuffling he chooses 4 cards out of 12 cards.

1st card - 12 ways
2nd card - 11 ways
3rd card - 10 ways
4th card - 9 ways

Total number of combinations = 12*11*10*9 / 4! = 495

The chance of Bill finding at least one pair of cards that have the same value
= 1 - chance of Bill finding no pair of cards that have same value

For a deck of 4 cards with all different cards

1st card - 12 ways
2nd card - 10 ways
3rd card - 8 ways
4th card - 6 ways

Number of combinations = 12*10*8*6/4! = 240

So the chance of Bill finding at least one pair of cards that have the same value = 1 - (240/495)
= [spoiler] 17/33 Answer C[/spoiler]
Join the discussion

by Brent@GMATPrepNow » Thu Jul 19, 2012 8:09 am
Ashishkapoor7 wrote:Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

8/33
62/165
17/33
103/165
25/33
I tackled this question by recognizing that P(at least one pair) = 1-P(no pairs)
First, the number of possible outcomes:
We have 12 cards and we select 4.
This is accomplished in 12C4 ways = 495

Now we count the number of ways to select 4 different values with no pairs. In other words, we wand 4 different card values.
First look at how many different cards we can select (WITHOUT considering the suits)
There are 6 card values and we want to select 4.
This is accomplished in 6C4 = 15 ways.
For each of these 15 values (e.g., 1,2,4,5) each card can be of either suit.
So, for example, in how many ways can we have card values of 1,2,4,5?
For the 1-card we have 2 suits from which to choose.
For the 2-card we have 2 suits from which to choose.
For the 4-card we have 2 suits from which to choose.
For the 5-card we have 2 suits from which to choose.
So, for the selection 1,2,4,5 there are 2x2x2x2-16 possibilities.
So, the number of ways to select 4 cards such that there are no pairs is 15x16=240
So, the probability that there are no pairs = 240/(11x5x9) = 48/99 = 16/33
So, P(at least one pair) = 1- 16/33 =17/33

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Brent@GMATPrepNow » Thu Jul 19, 2012 8:37 am
We can also solve this question using probability rules.

First, recognize that P(at least one pair) = 1 - P(no pairs)

P(no pairs) = P(select any 1st card AND select any non-matching card 2nd AND select any non-matching card 3rd AND select any non-matching card 4th)
= P(select any 1st card) x P(select any non-matching card 2nd) x P(select any non-matching card 3rd) x P(select any non-matching card 4th)
= 1 x 10/11 x 8/10 x 6/9
= 16/33

So, P(at least one pair) = 1 - 16/33
= 17/33
= C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by willrc » Thu Jul 19, 2012 11:16 am
We just need to decide how to avoid getting a pair:

Pick any card first (probability 1)
x
Avoid 1 out of 11 remaining cards (P=10/11)
x
Avoid 2 out of the 10 remaining cards (P=8/10)
x
Avoid 3 out of the 9 remaining cards (P=6/9).

1 x 10/11 x 8/10 x 6/9 = 16/33
1-16/33 = 17/33
London-based private tutor
www.gmatprolondon.com
Join the discussion